The width of a rectangle is 3 m less than twice its length. (By m I think it means meter, not a variable.) If the ratio of the length to the width is 7/5, find the dimensions of the rectangle.
Answer by jim_thompson5910(35256)

Length/Width = 7/5

L/(2L-3) = 7/5

5L = 7(2L-3)

5L = 7(2L-3)

5L = 14L-21

5L-14L = -21

-9L = -21

L = -21/(-9)

L = 7/3

W = 2L - 3

W = 2(7/3) - 3

W = 14/3 - 3

W = 14/3 - 9/3

W = 5/3

So the width is 5/3 m and the length is 7/3 m